Electric Force, Equilibrium Config of Charges Puntos:1 Find the charge Q that should be placed at the centre of the square of side 8.20E+0 cm, at the corners of which four identical charges +q = 5 C are placed so that the whole system is in equilibrium. Enviar Respuesta Tries os

Please explain me with steps. I dont understand how to solve it so well. Thank you
Answer #1

Solution :- Given data,

a = 8.20 cm, q = 5 C

the electric force on each q due to the other three q is

√[(kq2/a2)2 + (kq2/a2)2] + kq2/(a√2)2 = (√2 + 1/2)kq2/a2

so the net force on this q = (√2 + 1/2)kq2/a2 + kqQ/(a/√2)2 = 0

(√2 + 1/2)q= 2Q

Q = -q(√2 + 1/2)/2 = - 4.785 C